# Limit Computation Over Posets via Minimal Initial Functors

### v1

@@ 7. Proofs of Results on Limit Computation
In this section, we complete the proofs of our results on limit computation by proving Proposition˜3.26 and Corollary˜3.31 (i).
7.1. Proof of Proposition˜3.26
Recall the statement of Proposition˜3.26:
@@
We can compute a matrix representation of $H\circ i$ in time $O(|P|r^{\omega})$.
We prepare for the proof by recalling some results from elementary linear algebra, which we state without proof.
@@
For $\mathbbm{k}$-vector spaces $U\subseteq V$, we say a set of vectors $C$ in $V$ is a basis extension for $U\subseteq V$ if for some (or, equivalently, any) basis $B$ for $U$, $B\sqcup C$ is a basis for $V$.
@@
If $C$ is a basis extension for $U\subseteq V$, then the quotient map $V\to V/U$ sends $C$ to a basis $\bar{C}$ for $V/U$.
@@
Given a linear map $\alpha\colon V\to V^{\prime}$ with $\alpha(U)\subseteq U^{\prime}$, let $\bar{\alpha}\colon V/U\to V^{\prime}/U^{\prime}$ denote the induced map. Let $B,B^{\prime}$ be ordered bases for $U,U^{\prime}$, and let $C,C^{\prime}$ be ordered basis extensions for $U\subseteq V$, $U^{\prime}\subseteq V^{\prime}$. Then the matrix $\bar{\alpha}]^{\bar{C}^{\prime},\bar{C}}$ is the block of the matrix $\alpha]^{B^{\prime}\sqcup C^{\prime},B\sqcup C}$ consisting of columns indexed by $C$ and rows indexed by $C^{\prime}$.
@@
    Given a non-zero vector $c=(c_{1},\ldots,c_{l})\in\mathbbm{k}^{l}$, the pivot index of $c$ is the smallest index $k$ such that $c_{k}\neq 0$.
@@
    Given a finite dimensional vector space $V$ with ordered basis $A$ and $v\in V$, let $v]\in\mathbbm{k}^{|A|}$ be the representation of $v$ with respect to $A$. We say that an ordered set of vectors $C=\{c_{1},\ldots,c_{l}\}$ in $V$ is in echelon form (with respect to $A$) if the pivot indices of the $c_{k}]$ are strictly increasing with $k$.
Proof of Proposition˜3.26.
Recalling Definition˜3.25, suppose that the given matrices $[f]$ and $[g]$ represent $f\colon X\to Y$ and $g\colon Y\to Z$ with respect to ordered bases $A^{X},A^{Y},A^{Z}$ for $X,Y,Z$. For each $p\in P$, these bases induce ordered bases $A^{X}_{p},A^{Y}_{p},A^{Z}_{p}$ for $X_{p},Y_{p},Z_{p}$. The matrix representation $f_{p}]$ of $f_{p}$ with respect to the induced bases is obtained from $[f]$ by taking the submatrix consisting of all rows and columns with labels $\leq p$, and similarly for $g_{p}]$. Since we assume that pairs of elements of $Q$ can be compared in constant time, deciding whether a column or row of $[f]$ belongs to $f_{p}]$ also takes constant time.
@@
We next explain the computation of $B_{p}$, $C_{p}$, and $D_{p}$ in detail. To begin, we compute an ordered basis $A^{\ker g}_{p}$ of $\ker g_{p}$ in echelon form (with respect to $A^{Y}_{p}$). We do this by applying Gaussian elimination with backsolve to $g_{p}]$. Writing $A^{Y}_{p}=\{a_{1},\ldots,a_{l}\}$, we take
@@
For each relation $p\leq q$ in $P$, we must compute the matrix $H_{pq}]^{\bar{C}_{q},\bar{C}_{p}}$. To do this, first we form the matrix
$Y_{pq}]^{A^{Y}_{q},A^{Y}_{p}}=\begin{pmatrix}I\\ 0\end{pmatrix}$.
We then multiply this matrix by change-of-basis matrices on both the left and right to obtain the matrix
@@
It remains to show that this algorithm runs in time $O(|P|r^{\omega})$. For each element or relation of $P$, the algorithm performs Gaussian elimination with backsolve a constant number of times and also performs a constant number of matrix multiplications. Each matrix considered has dimension $O(r)\times O(r)$, so each instance of Gaussian elimination or matrix multiplication requires time $O(r^{\omega})$; see Proposition˜3.21. Since such operations dominate the cost of the algorithm, it follows that our algorithm runs in time $O(|P|r^{\omega})$. ∎
7.2. Proof of Corollary˜3.31 (i)
Let us recall the statement of Corollary˜3.31 (i), i.e., the case $d=2$ of the corollary:
@@
Given an $(\mathbb{N}^{2},r)$-complex with homology $H$ and the upset presentation of an interval $i\colon Q\hookrightarrow\mathbb{N}^{2}$ with $s=\|Q\|$, we can compute $\lim(H\circ i)$ in time $O(s\log s+r^{3})$.
@@
A standard result in quiver representation theory [gabriel1972unzerlegbare, carlsson2010zigzag] says that for $P$ a finite zigzag poset (see Example˜3.4 and a functor $G\colon P\to\mathbf{Vec}$, there is a unique multiset $\mathcal{B}(G)$ of intervals in $P$ such that
(7.1) $G\cong\bigoplus_{I\in\mathcal{B}(G)}\mathbbm{k}^{I}$
where the $\mathbbm{k}^{I}$ are the interval modules of Definition˜2.14. If all vector spaces of $G$ have finite dimension, then $\mathcal{B}(G)$ is finite, so the direct sum decomposition of $G$ above is in fact a direct product. Since small limits and categorical products commute [riehl2017category, Theorem 3.8.1], we have
$\lim G\cong\bigoplus_{I\in\mathcal{B}(G)}\lim\mathbbm{k}^{I}.$
Moreover, it is easily checked that for each interval $I\subseteq P$, we have
$\lim\mathbbm{k}^{I}=\begin{cases}\mathbbm{k}&\textup{ if $I$ is a downset},\\ 0&\textup{otherwise.}\end{cases}$
It follows that we can compute $\lim G$ by computing a decomposition of $G$ into indecomposable modules $\mathbbm{k}^{I}$ as in Eq.˜7.1. However, an important subtlety here is that, to compute a limit cone for $G$ (and not only for some functor naturally isomorphic to $G$), it is not enough to compute $\mathcal{B}(G)$; we also need an explicit isomorphism
(7.2) $\bigoplus_{I\in\mathcal{B}(G)}\mathbbm{k}^{I}\xrightarrow{\cong}G.$
We now apply these ideas to our setting. By Example˜3.4, $Q$ has a unique initial scaffold $j\colon P\hookrightarrow Q$, where $P$ is a finite zigzag poset. Let $G$ be the restriction of $H$ to $P$, i.e., $G=H\circ i\circ j$. We can compute $\mathcal{B}(G)$ and the isomorphism of Eq.˜7.2 in time $O(r^{3})$ by applying the zigzag persistence algorithm of [DTM25] to the restriction of the chain complex $\operatorname{\mathcal{C}}_{\mathbb{N}^{2}}^{r}$ to $P$. By Theorem˜3.17 (i), we can compute $P$ in time $O(s\log s)$. These two computations dominate the cost of computing $\lim(G)$. Hence, this approach computes $\lim G$, and thus $\lim(H\circ i)$, in total time $O(s\log s+r^{3})$. ∎
@@
If we wish to only compute $\dim(\lim(G\circ i))$ or the limit cone of a functor naturally isomorphic to $G\circ i$, then the complexity bound of Corollary˜3.31 (i) can be improved $O(s\log s+r^{\omega})$ by using either the zigzag persistence algorithm of [milosavljevic2011zigzag] or the more practical algorithm of [DH22].
@@ 8. Proofs of Results on Generalized Rank Computation
In this section, we prove Corollary˜3.36 and Corollary˜3.37, our bounds on the cost of generalized rank computation. We use the following notation:
@@
We first prove Corollary˜3.36. Recall the statement:
@@
Let $Q$ be a finite, connected poset with $s$ extrema and let $^{\mathcal{I}}, PP^{\mathcal{F}}$ denote initial and final scaffolds of $Q$. Given the Hasse diagram $(V,E)$ of $Q$ and a $(Q,r)$-complex with homology $H$, we can compute $\operatorname{grank}(H)$ in time
$O(|V||E|+(|P^{\mathcal{I}}|+|P^{\mathcal{F}}|)s^{\omega-1}r^{\omega}).$
@@
To compute $\operatorname{grank}(H)$, we first compute initial and final scaffolds $j^{\mathcal{I}}\colon P^{\mathcal{I}}\hookrightarrow Q$ and $j^{\mathcal{F}}\colon P^{\mathcal{F}}\hookrightarrow Q$ via Theorem˜3.13. Choosing $m\leq w$ as in ˜8.1 and letting $j\colon P\hookrightarrow Q$ be as defined there, we compute a matrix representation of $H\circ j$ via Propositions˜3.26 and 3.29, which, in particular, yields a basis $B_{p}$ of $H_{p}$ for each $p\in P$. This restricts to matrix representations of $H\circ j^{\mathcal{I}}$ and $H\circ j^{\mathcal{F}}$, which we use to compute $\lim(H\circ j^{\mathcal{I}})$ and $\operatorname{colim}(H\circ j^{\mathcal{F}})$ as in Corollary˜3.30.
Let $\gamma_{m}\colon\lim H\to H_{m}$ and $\gamma^{\prime}_{w}\colon H_{w}\to\operatorname{colim}H$ be the respective cone and cocone maps. Observe that
$\operatorname{grank}H=\operatorname{rank}(\gamma^{\prime}_{w}\circ H_{mw}\circ\gamma_{m}).$
Since $m$ is a minimum, a matrix representation $\gamma_{m}]$ of $\gamma_{m}$ with respect to the basis $B_{m}$ is given by coordinate projection of the computed presection basis of $\lim H$; see ˜3.19 (i). Dually, coordinate projection also yields a matrix representation $\gamma^{\prime}_{w}]$ of $\gamma^{\prime}_{w}$ with respect to the basis $B_{w}$. The matrix representation of $H\circ j$ furnishes a matrix representation of $H_{mw}]$ with respect to the bases $B_{m}$ and $B_{w}$. To compute $\operatorname{grank}H$, we compute the matrix product $\gamma^{\prime}_{w}][H_{mw}][\gamma_{m}]$ and perform Gaussian elimination on the result.
The cost of these computations is dominated by the cost of computing $\lim H$ and $\operatorname{colim}H$, which is given by Corollary˜3.30. The result follows. ∎
We next turn to the proof of Corollary˜3.37. Recall the statement:
@@
Given an $(\mathbb{N}^{d},r)$-complex with homology $H$ and the $s$ extrema of a finite interval $i\colon Q\hookrightarrow\mathbb{N}^{d}$, we can compute $\operatorname{grank}(H\circ i)$ in time
@@
    $O(r^{\omega}s^{\omega+1})$ for $d>3$.
Proof of Corollary˜3.37 (ii,iii).
The proof is the same as the proof of Corollary˜3.36, using Theorems˜3.17, 3.31 and 3.32 in place of Theorems˜3.13 and 3.30. ∎
In the setting of Corollary˜3.37 (i), the proof of Corollary˜3.36, together with Corollary˜3.31 (i), yields the bound $O(s\log s+r^{3})$ for generalized rank computation over a finite interval in $\mathbb{N}^{2}$, but does not yield the stronger bound $O(s\log s+r^{\omega})$ given by Corollary˜3.37 (i). The term $r^{3}$ arises from the cost of (co)limit computation via Corollary˜3.31 (i); see Remark˜7.5. We therefore need a different argument to prove Corollary˜3.37 (i). Our proof hinges on following lemma.
@@
We will only need the case of Proposition˜8.2 where $Q$ is an interval in $\mathbb{N}^{2}$, but it is no more difficult to prove the general result. As noted in Remark˜3.7, a variant of the proposition was obtained independently in [brustle2025generalized].
@@
By the functoriality of limits with respect to the index category, this isomorphism factors as
(8.1) $\lim G\hookrightarrow\lim(G\circ j)\twoheadrightarrow\lim(G\circ j^{\mathcal{I}}),$
where the first map is an injection, the second map is a surjection, and both maps commute with the relevant cone maps.
Dually, the map $j^{\mathcal{F}}$ induces an isomorphism $\operatorname{colim}(G\circ j^{\mathcal{F}})\to\operatorname{colim}G$ and by functoriality, this map factors as
(8.2) $\operatorname{colim}(G\circ j^{\mathcal{F}})\hookrightarrow\operatorname{colim}(G\circ j)\twoheadrightarrow\operatorname{colim}G,$
where again, the first map is an injection and the second map is an surjection. Let $\alpha$ denote the composite map
@@
Note that $\operatorname{grank}(G\circ j)=\operatorname{rank}(\alpha)$. Moreover, since $j^{\mathcal{I}}$ and $j^{\mathcal{F}}$ are initial and final, respectively, we also have $\operatorname{grank}(G)=\operatorname{rank}\beta$. The map $\alpha$ factors as
$\lim(G\circ j)\twoheadrightarrow\underbrace{\lim(G\circ j^{\mathcal{I}})\to G_{m}\to G_{w}\to\operatorname{colim}(G\circ j^{\mathcal{F}})}_{\beta}\hookrightarrow\operatorname{colim}(G\circ j),$
where the first factor is the surjection from Eq.˜8.1 and the last factor is the injection from Eq.˜8.2.
Observe that for any linear maps $f\colon V^{\prime}\to V$, $g\colon V\to W$, and $h\colon W\to W^{\prime}$ with $f$ a surjection and $h$ an injection, we have $\operatorname{rank}(h\circ g\circ f)=\operatorname{rank}g$. Applying this to the maps above yields
@@
Proof of Corollary˜3.37 (i).
We use an argument similar to the proof of Theorem˜3.35 in [DKM22], but framed in our general formalism of initial and final scaffolds. We compute initial and final scaffolds $P^{\mathcal{I}}$ and $P^{\mathcal{F}}$ of $Q$ in time $O(s\log s)$ using Theorem˜3.17 (i). Recall from Example˜3.4 that $P^{\mathcal{I}}$ is a zigzag poset; by duality, the same is true for $P^{\mathcal{F}}$. Let $m$ and $w$ be the leftmost minimum and maximum of $Q$, respectively. Then $m\in P^{\mathcal{I}}$, $w\in P^{\mathcal{F}}$, and $m\leq w$. Let $P$ and $j$ be as in ˜8.1 and let $G=H\circ i$. Then $P$ is a zigzag poset and the barcode of the zigzag persistence module $G\circ j$ can be computed in time $O(s\log s+r^{\omega})$ via the algorithm of either [milosavljevic2011zigzag] or [DH22]. The number of copies of the interval $P$ in the barcode is the number of interval summands of $G\circ j$ with support $P$, which in turn is $\operatorname{grank}(G\circ j)$ by [chambers2018persistent, Corollary 3.2]. We have $\operatorname{grank}(G\circ j)=\operatorname{grank}(G)=\operatorname{grank}(H\circ i)$ by Proposition˜8.2. The result follows. ∎

### v2

@@ 7. Proofs of Results on Limit Computation
In this section, we complete the proofs of our results on limit computation by proving Proposition˜3.25 and Corollary˜3.30 (i).
7.1. Proof of Proposition˜3.25
Recall the statement of Proposition˜3.25:
@@
Then we can compute a matrix representation of $H\circ i$ in time $O(|P|r^{\omega})$.
We prepare for the proof by recalling some elementary linear algebra results, which we state without proof.
@@
For $\mathbbm{k}$-vector spaces $U\subseteq V$, we say a set of vectors $C$ in $V$ is a basis extension of $U\subseteq V$ if $B\sqcup C$ is a basis of $V$ whenever $B$ is a basis of $U$.
@@
If $C$ is a basis extension of $U\subseteq V$, then the quotient map $V\to V/U$ sends $C$ to a basis $\bar{C}$ of $V/U$.
@@
Given a linear map $\alpha\colon V\to V^{\prime}$ with $\alpha(U)\subseteq U^{\prime}$, let $\bar{\alpha}\colon V/U\to V^{\prime}/U^{\prime}$ denote the induced map. Let $B,B^{\prime}$ be ordered bases for $U,U^{\prime}$, and let $C,C^{\prime}$ be ordered basis extensions of $U\subseteq V$, $U^{\prime}\subseteq V^{\prime}$. Then the matrix $\bar{\alpha}]^{\bar{C}^{\prime},\bar{C}}$ is the block of the matrix $\alpha]^{B^{\prime}\sqcup C^{\prime},B\sqcup C}$ consisting of rows indexed by $C^{\prime}$ and columns indexed by $C$.
@@
    Given a non-zero vector $c=(c_{1},\ldots,c_{l})\in\mathbbm{k}^{l}$, the pivot index of $c$ is the smallest index $j$ such that $c_{j}\neq 0$.
@@
    Given a finite dimensional vector space $V$ with ordered basis $A$ and $v\in V$, let $v]\in\mathbbm{k}^{|A|}$ be the representation of $v$ with respect to $A$. We say that an ordered set of vectors $C=(v_{1},\ldots,v_{l})$ in $V$ is in echelon form (with respect to $A$) if the pivot indices of the vectors $v_{k}]$ are strictly increasing with $k$.
Proof of Proposition˜3.25.
Recalling Definition˜3.24, suppose that the given matrices $[f]$ and $[g]$ represent $f\colon X\to Y$ and $g\colon Y\to Z$ with respect to ordered bases $A^{X},A^{Y},A^{Z}$ for $X,Y,Z$. For each $p\in P$, these bases induce ordered bases $A^{X}_{p},A^{Y}_{p},A^{Z}_{p}$ for $X_{p},Y_{p},Z_{p}$. The matrix representation $f_{p}]$ of $f_{p}$ with respect to the induced bases is obtained from $[f]$ by taking the submatrix consisting of all rows and columns with labels $\leq p$. Since we assume that pairs of elements of $Q$ can be compared in constant time, deciding whether a column or row of $[f]$ belongs to $f_{p}]$ also takes constant time. The same is true for $g_{p}]$.
@@
We now explain this computation in detail. To begin, we compute an ordered basis $A^{\ker g}_{p}$ of $\ker g_{p}$ in echelon form (with respect to $A^{Y}_{p}$). We do this by applying Gaussian elimination with backsolve to $g_{p}]$. Writing $A^{Y}_{p}=\{a_{1},\ldots,a_{l}\}$, we take
@@
For each relation $p\leq q$ in $P$, we must compute the matrix $H_{pq}]^{\bar{C}_{q},\bar{C}_{p}}$. To do this, first we form the matrix $Y_{pq}]^{A^{Y}_{q},A^{Y}_{p}}$. Note that this is the binary matrix whose entry at index $(y,z)$ is 1 if and only if the $y^{\mathrm{th}}$ element of $A^{Y}_{q}$ and the $z^{\mathrm{th}}$ element of $A^{Y}_{p}$ are induced by the same element of $A^{Y}$. We then multiply this matrix by change-of-basis matrices on both the left and right to obtain the matrix
@@
It remains to show that this algorithm runs in time $O(|P|r^{\omega})$. For each element or relation of $P$, the algorithm performs Gaussian elimination with backsolve a constant number of times and also performs a constant number of matrix multiplications. Each matrix considered has dimension $O(r)\times O(r)$, so each instance of Gaussian elimination or matrix multiplication requires time $O(r^{\omega})$; see Proposition˜3.20. Since such operations dominate the cost of the algorithm, it follows that our algorithm runs in time $O(|P|r^{\omega})$. ∎
7.2. Proof of Corollary˜3.30 (i)
Let us recall the statement of Corollary˜3.30 (i):
@@
Given an $(\mathbb{N}^{2},r)$-complex with homology $H$ and the upset presentation of an interval $i\colon Q\hookrightarrow\mathbb{N}^{2}$ with $\|Q\|=n$, we can compute $\lim(H\circ i)$ in time $O(n\log n+r^{3})$.
Our proof will use the following decomposition result from classical quiver representation theory, due to Gabriel.
Theorem 7.5 ([24, 9]).
Given a zigzag poset $P$ (see Definition˜2.5) and a pointwise-dimensional functor $G\colon P\to\mathbf{Vec}$, there is a unique finite multiset $\mathcal{B}(G)$ of intervals in $P$ such that
(7.1) $G\cong\bigoplus_{I\in\mathcal{B}(G)}\mathbbm{k}^{I},$
where the $\mathbbm{k}^{I}$ are the interval modules of Definition˜2.18.
Proposition 7.6.
In the setting of Theorem˜7.5, we have
(7.2) $\lim G\cong\bigoplus_{I\in\mathcal{B}(G)}\lim\mathbbm{k}^{I},$
where
$\lim\mathbbm{k}^{I}=\begin{cases}\mathbbm{k}&\textup{ if $I$ is a downset of $P$},\\ 0&\textup{otherwise.}\end{cases}$
@@
Since $\mathcal{B}(G)$ is finite, the direct sum of Eq.˜7.1 is in fact a direct product. Thus, since small limits and categorical products commute [39, Theorem 3.8.1], Eq.˜7.2 holds. The formula for $\lim\mathbbm{k}^{I}$ follows from Proposition˜2.10. ∎
Example 7.7.
Let $P$ be the zigzag poset from Example˜2.6, whose Hasse diagram is
v{v}w{w}x{x}y{y}z.{z.}
Note that $\{v,w,x\}$ is a downset of $P$, but $\{v,w,x,y\}$ is not. Therefore, the following two diagrams have limits $\mathbbm{k}$ and 0, respectively:
[…]
By Proposition˜7.6, to compute $\lim G$ in the setting of Theorem˜7.5, it suffices to compute a decomposition of $G$ as in Eq.˜7.1. However, an important subtlety here is that, to compute a limit cone for $G$ (and not only for some functor naturally isomorphic to $G$), it is not enough to compute $\mathcal{B}(G)$; we also need an explicit choice of the isomorphism of Eq.˜7.1.
Proof of Corollary˜3.30 (i).
By Example˜3.4, $Q$ has a unique initial scaffold $j\colon P\hookrightarrow Q$, where $P$ is a finite zigzag poset. Let $G$ be the restriction of $H$ to $P$, i.e., $G=H\circ i\circ j$. We can compute $\mathcal{B}(G)$ and the isomorphism of Eq.˜7.1 in time $O(r^{3})$ by applying the zigzag persistence algorithm of Dey, Hao, and Morozov [18] to the restriction of the input $(\mathbb{N}^{2},r)$-complex to $P$. By Theorem˜3.16 (i), we can compute $P$ in time $O(n\log n)$. These two computations dominate the cost of computing $\lim(G)$. Hence, this approach computes $\lim G$, and thus $\lim(H\circ i)$, in total time $O(n\log n+r^{3})$. ∎
@@
If we wish to only compute $\dim(\lim(G\circ i))$ or the limit cone of a functor naturally isomorphic to $G\circ i$, then the complexity bound of Corollary˜3.30 (i) can be improved to $O(n\log n+r^{\omega})$ by using either the zigzag persistence algorithm of Milosavljević, Morozov, and P. Skraba [36] or the more practical algorithm of Dey and Hao [19]. However, these algorithms do not provide the isomorphism of Eq.˜7.1, which we need to compute the cone maps of $\lim(G\circ i)$.
@@ 8. Proofs of Results on Generalized Rank Computation
In this section, we prove Corollary˜3.35 and Corollary˜3.36, our bounds on the cost of generalized rank computation. We use the following notation:
@@
We first prove Corollary˜3.35. Recall the statement:
@@
Let $Q$ be a finite, connected poset with $n$ extrema and let $^{\mathcal{I}}, PP^{\mathcal{F}}$ be initial and final scaffolds of $Q$. Given the Hasse diagram $(V,E)$ of $Q$ and a $(Q,r)$-complex with homology $H$, we can compute $\operatorname{grank}(H)$ in time
$O(|V||E|+(|P^{\mathcal{I}}|+|P^{\mathcal{F}}|)n^{\omega-1}r^{\omega}).$
@@
To compute $\operatorname{grank}(H)$, we first compute initial and final scaffolds $j^{\mathcal{I}}\colon P^{\mathcal{I}}\hookrightarrow Q$ and $j^{\mathcal{F}}\colon P^{\mathcal{F}}\hookrightarrow Q$ via Theorem˜3.13. Choosing $m\leq w$ as in ˜8.1 and letting $j\colon P\hookrightarrow Q$ be as defined there, we compute a matrix representation of $H\circ j$ via Propositions˜3.25 and 3.28, which, in particular, yields a basis $B_{p}$ of $H_{p}$ for each $p\in P$. This matrix representation restricts to matrix representations of $H\circ j^{\mathcal{I}}$ and $H\circ j^{\mathcal{F}}$, which we use to compute $\lim(H)$ and $\operatorname{colim}(H)$ as we did in Corollary˜3.29.
Let $\delta_{m}\colon\lim H\to H_{m}$ and $\delta^{\prime}_{w}\colon H_{w}\to\operatorname{colim}H$ be the respective cone and cocone maps. Observe that
$\operatorname{grank}(H)=\operatorname{rank}(\delta^{\prime}_{w}\circ H_{mw}\circ\delta_{m}).$
Since $m$ is a minimum, a matrix representation $\delta_{m}]$ of $\delta_{m}$ with respect to the basis $B_{m}$ is given by coordinate projection of the computed presection basis of $\lim H$; see ˜3.18 (i). Dually, coordinate projection also yields a matrix representation $\delta^{\prime}_{w}]$ of $\delta^{\prime}_{w}$ with respect to the basis $B_{w}$. The matrix representation of $H\circ j$ furnishes a matrix representation of $H_{mw}]$ with respect to the bases $B_{m}$ and $B_{w}$. To compute $\operatorname{grank}(H)$, we compute the matrix product $\delta^{\prime}_{w}][H_{mw}][\delta_{m}]$ and perform Gaussian elimination on this.
The cost of these computations is dominated by the cost of computing $\lim H$ and $\operatorname{colim}H$, which is given by Corollary˜3.29. The result follows. ∎
We next turn to the proof of Corollary˜3.36. Recall the statement:
@@
Given an $(\mathbb{N}^{d},r)$-complex with homology $H$ and the $n$ extrema of a finite interval $i\colon Q\hookrightarrow\mathbb{N}^{d}$, we can compute $\operatorname{grank}(H\circ i)$ in time
@@
    $O(n^{4}+n^{\omega+1}r^{\omega})$ for $d>3$.
Proof of Corollary˜3.36 (ii,iii).
The proof is the same as the proof of Corollary˜3.35, using Theorem˜3.16 in place of Theorem˜3.13, as well as Corollaries˜3.30 and 3.31 in place of Corollary˜3.29. ∎
In the setting of Corollary˜3.36 (i), the proof of Corollary˜3.35, together with Corollary˜3.30 (i), yields the bound $O(n\log n+r^{3})$ for generalized rank computation over a finite interval in $\mathbb{N}^{2}$, but does not yield the stronger bound $O(n\log n+r^{\omega})$ given by Corollary˜3.36 (i). The term $r^{3}$ arises from the cost of (co)limit computation via Corollary˜3.30 (i); see Remark˜7.8. We therefore need a different argument to prove Corollary˜3.36 (i). Our proof hinges on the following lemma.
@@
We will only need the case of Proposition˜8.2 where $Q$ is an interval in $\mathbb{N}^{2}$, but it is no more difficult to prove the general result. As noted in Remark˜3.7, a variant of the proposition was obtained independently in [8].
@@
By the functoriality of limits with respect to the index category (see Section˜2.2), this isomorphism factors as
$\lim G$$\lim(G\circ j)$$\lim(G\circ j^{\mathcal{I}})$$\scriptstyle{\rho}$
where the first map is an injection and the second map is a surjection. Dually, the map $j^{\mathcal{F}}$ induces an isomorphism $\operatorname{colim}(G\circ j^{\mathcal{F}})\to\operatorname{colim}G$ factoring as
$\operatorname{colim}(G\circ j^{\mathcal{F}})$$\operatorname{colim}(G\circ j)$$\operatorname{colim}G$$\scriptstyle{\iota}$
where again, the first map is an injection and the second map is a surjection. Let $\alpha$ denote the composite map
@@
Then following diagram commutes:
$\lim(G\circ j)$$\operatorname{colim}(G\circ j)$$G_{m}$$G_{w}$$\lim(G\circ j^{\mathcal{I}})$$\operatorname{colim}(G\circ j^{\mathcal{F}})$$\scriptstyle{\rho}$$\scriptstyle{\alpha}$$\scriptstyle{\beta}$$\scriptstyle{\iota}$
Thus, $\alpha=\iota\circ\beta\circ\rho$. Since $\rho$ is a surjection and $\iota$ is an injection, we have $\operatorname{rank}(\alpha)=\operatorname{rank}(\beta)$. By definition, we have $\operatorname{grank}(G\circ j)=\operatorname{rank}(\alpha)$. Moreover, since $j^{\mathcal{I}}$ and $j^{\mathcal{F}}$ are initial and final, respectively, Proposition˜2.14 and its dual imply that $\operatorname{grank}(G)=\operatorname{rank}(\beta)$. Therefore,
@@
Proof of Corollary˜3.36 (i).
We use an argument similar to the proof of Theorem˜3.34 in [16], but framed in our general formalism of initial and final scaffolds. We compute initial and final scaffolds $P^{\mathcal{I}}$ and $P^{\mathcal{F}}$ of $Q$ in time $O(n\log n)$ using Theorem˜3.16 (i). Recall from Example˜3.4 that $P^{\mathcal{I}}$ is a zigzag poset; by duality, the same is true for $P^{\mathcal{F}}$. Let $m$ and $w$ be the leftmost minimum and maximum of $Q$, respectively. Then $m\in P^{\mathcal{I}}$, $w\in P^{\mathcal{F}}$, and $m\leq w$. Let $P$ and $j$ be as in ˜8.1, and let $G=H\circ i$. Then $P$ is a zigzag poset and the barcode of the zigzag module $G\circ j$ can be computed in time $O(n\log n+r^{\omega})$ via the algorithm of either [36] or [19]. The number of copies of the interval $P$ in the barcode is the number of interval summands of $G\circ j$ with support $P$, which in turn is $\operatorname{grank}(G\circ j)$ by [13, Corollary 3.2]. We have $\operatorname{grank}(G\circ j)=\operatorname{grank}(G)=\operatorname{grank}(H\circ i)$ by Proposition˜8.2. The result follows. ∎